For each of the following sets of volume/temperature data, calculate the missing quantity after the change is made. Assume that the pressure and the amount of gas remain the same.
(a) V = 2.06 L at 28°C; V = 3.41 L at WebAssign will check your answer for the correct number of significant figures. °C

(b) V = 105 mL at 221 K; V = WebAssign will check your answer for the correct number of significant figures. mL at 376 K

(c) V = 48.8 mL at 40.°C; V = WebAssign will check your answer for the correct number of significant figures. mL at 369 K
Answer #1

According to the Charle's law

For a gas , temperature and volume are directly proportional . . Here pressure and amount of gas is remain same .

V \alpha T

V = KT

Where V = volume of the gas

K = proportionality constant

T = temperature in kelvin

Or \frac{V}{T}= K

Applying Charle's Law in different temperature and volume for a gas

SK TV and   \frac{V_{2}}{T_{2}}= K

Hence

  V1 Ti V2 Tə

a) here V 1 = 2.06 L

T​​​​​1 = 28°C = (28+273.15)K = 301.15K

V 2 = 3.41 L

Using Charle's law

  V1 Ti V2 Tə

3.412 2.062 301.15K T2

T​​​​​​2 = 498.50K

Hence the temperature is (498.50-273.15)K = 225.35 °C

b) Here V​​​​​​1 = 105 mL

T​​​​​​1 = 221 K

V​​​​​​2 = unknown

T​​​​​​2 = 376 K

Using Charle's law

  V1 Ti V2 Tə

105mL 221K V2 376K

V2 = 105mL x 376K 221K

V​​​​​2 = 178.64 mL

Hence the volume is 178.64 mL

C) here   

V​​​​​​1 = 48.8 mL

T​​​​​​1 = 40°C= (40+273.15) K = 313.15 K

V​​​​​​2 = unknown

T​​​​​​2 = 369 K

Using Charle's law

V1 Ti V2 Tə

  48.8mL 313.15K V2 369K

  V2 =- 48.8mL x 369K 313.15K

V​​​​​​2 = 57.50 mL

Hence the volume is 57.50 mL .